OpenStudy (anonymous):

using complete sentences, explain how you would use quadratic formula to solve x^2 +5x=-2. why is the quadratic formula the best method to use ? (7 points)

OpenStudy (anonymous):

doesnt say how to solve it

OpenStudy (anonymous):

I dont need that

OpenStudy (nurali):

x^2 + 5x +2 = 0 Formula: x=-b+-Sqrt(b^2-4ac)/2a your coefficients are a = 1 b = 5 c = 2 then, use the formula and you get x1 = [ -5 + sqrt( 25 - 8) ] / 2 = -0.4384 x2 = [ -5 - sqrt( 25 - 8) ] / 2 = -4.5616

OpenStudy (anonymous):

if its decimals its wrong my teacher said

OpenStudy (anonymous):

@Nurali